OpenStudy (anonymous):

4x^2+32x-(-48)=0 Helppppp

OpenStudy (anonymous):

4x^2+32x-(-48)=0 negative and negative gives you positive 4x^2+32x+48=0 Solve it like a quadratic equation

OpenStudy (anonymous):

how do you do 4(x+6)(x+2) ? do you distribute the 4 to both x+6 and x+2 or just x+6?

OpenStudy (anonymous):

both of them since you factorize out from both. One way to check is to multiply the 4 back into brackets and see if you get back your original equation

OpenStudy (anonymous):

when i did it i only got \[4x ^{2}+32\]

OpenStudy (anonymous):

i didnt end up with the -(-48)

OpenStudy (anonymous):

How did you get 4x^2+32?

OpenStudy (anonymous):

4(x+6)(x+2)

OpenStudy (anonymous):

unless I'm doing it wrong

OpenStudy (anonymous):

you multiply the four into the first bracket 4x+24(x+2) Multiply each term with both the terms inside bracket 4x*x+4x*2+24*x+24*2 =4x^2+8x+24x+48 =4x^2 +32x + 48 =4x^2 + 32x -(-48)

OpenStudy (anonymous):

ohhhhhh okay , i see!

OpenStudy (anonymous):

another question , if you dont mind